I have this question 2x...different answers to pick from...please help me out??
Find the slope of the line passing through the points (–2, 7) and (8, –1).
1st set of answers...
a.)5/4
b.)-5/4
c.)4/5
d.)-4/5
2nd set of answers...
a.)0
b.)1
c.)-4
d.) Undefined
how is this possible and can it broken down into steps?

Answer by checkley75(3666)

SOMETIMES THE QUESTION ISN'T CLEAR BECAUSE THE USE OF () ARE NOT USED OFTEN ENOUGH.
EXAMPLE 1/2X CAN BE INTERPRETED AS (1/2)X OR 1/(2X) THESE ARE 2 TOTALLY DIFFERENT TERMS.
MATH IS A PRECISE SCIENCE & THER IS ONLY 1 CORRECT ANSWER HOWEVER THE ANSWER CAN BE PRESENTED IN DIFFERENT FORMS;
EXAMPLE 20%=.2=1/5 ALL ARE THE SAME ANSWER - JUST IN DIFFERENT FORMS.
SLOPE=(Y2-Y1)/(X2-X1)
SLOPE=(-1-7)/(8+2)
SLOPE=-8/10
SLOPE=-4/5 OR ANSWER D).